OpenStudy (anonymous):

A number cube is rolled 360 times and the results as recorded as followed: 54 ones 60 twos 66 threes 71 fours 35 fives and 74 sixes what is the experimental probrobility you will roll a two or a three

OpenStudy (anonymous):

Chance you will roll a two = 60 / 360 = 0.16666 ≈ 0.17 ≈ 17% Chance you will roll a three = 66 / 360 = 0.183333 ≈ 0.18 ≈ 18%

OpenStudy (anonymous):

To give out medals click 'Best Response'.

OpenStudy (anonymous):

@angelrush